# The Hidden Costs of Translation Accuracy: Distillation, Quantization, and Environmental Impact

### v1

@@ Abstract
The rapid expansion of large language models (LLMs) has heightened concerns about their computational and environmental costs. This study investigates the trade-offs between translation quality and efficiency by comparing full-scale, distilled, and quantized models using machine translation as a case study. We evaluated performance on the Flores+ benchmark and through human judgments of conversational translations in French, Hindi, and Kannada. Our analysis of carbon emissions per evaluation run revealed that the full 3.3B fp32 model, while achieving the highest BLEU scores, incurred the largest environmental footprint (about 0.007–0.008 kg per run). The distilled models achieved an inference of up to 4.5 faster than the full 3.3B model, with only minimal reductions in BLEU scores. Human evaluations also showed that even aggressive quantization (INT4) preserved high levels of accuracy and fluency, with differences between models generally minor. These findings demonstrate that model compression strategies can substantially reduce computational demands and environmental impact while maintaining competitive translation quality, though trade-offs are more pronounced in low-resource settings. We argue for evaluation frameworks that integrate efficiency and sustainability alongside objective metrics as central dimensions of progress in NLP.
@@ 4.1 Benchmark Evaluation (Flores+)
Despite these reductions in BLEU scores, the distilled model achieved substantial efficiency gains. Inference was 3.5–4.5 faster across all three language pairs, with average sentence latency dropping from 0.187s to 0.041s for English-French and from 0.146s to 0.042s for English-Kannada. Carbon emissions per evaluation run decreased by more than half, from 0.0075-0.0079 kg to 0.0025-0.0029 kg . These results demonstrate that distillation significantly reduced computational and environmental costs, albeit with modest quality degradation.
| BLEU Score | Inference Time | Carbon Emissions | ||||
| Language Pair | Full | Distilled | Full | Distilled | Full | Distilled |
| English-French | 43.99 | 42.10 | 0.187 | 0.041 | 0.0075 | 0.0026 |
| English-Kannada | 15.31 | 11.80 | 0.146 | 0.042 | 0.0079 | 0.0029 |
| English-Hindi | 28.67 | 27.42 | 0.134 | 0.037 | 0.0069 | 0.0025 |
@@ 5 Conclusion
The distilled models achieved up to 4.5× faster inference and reduced carbon emissions by more than half compared with the full model, while incurring only modest losses in BLEU scores. Performance degradation was most pronounced for low-resource Kannada, whereas high-resource French remained largely stable.
@@
These results demonstrate that distillation and quantization can substantially reduce computational demands and environmental costs without severely compromising translation quality. However, the sharper trade-offs observed in low-resource languages highlight the need for careful evaluation before deployment in linguistically diverse settings.
Our analysis was limited to inference-time efficiency; the training phase, which contributes significantly to the total carbon footprint of large models, remains an important direction for further research. Moreover, the study focused on three languages; expanding to a broader set would provide a more comprehensive understanding of multilingual trade-offs.
Overall, these findings support the case for evaluation frameworks that integrate efficiency and sustainability alongside accuracy. Future work should extend this analysis to additional languages and tasks and incorporate lifecycle assessments of training, deployment, and model updates, thereby advancing the agenda of responsible and environmentally conscious NLP.

### v2

@@ Abstract
The rapid expansion of large language models (LLMs) has heightened concerns about their computational and environmental costs. This study investigates the trade-offs between translation quality and efficiency by comparing full-scale, distilled, and quantized models using machine translation as a case study. We evaluated performance on the Flores+ benchmark and through human judgments of conversational translations in French, Hindi, and Kannada. Our analysis revealed that the full 3.3B FP32 model, while achieving the highest BLEU scores, incurred the largest environmental footprint ( 0.007–0.008 kg per run). The distilled 600M FP32 model reduced inference time by 71–78% and carbon emissions by 63–65% compared with the full model, with only minimal reductions in BLEU scores. Human evaluations further showed that even aggressive quantization (INT4) preserved high levels of accuracy and fluency, with differences between models generally minor. These findings demonstrate that model compression strategies can substantially reduce computational demands and environmental impact while maintaining competitive translation quality, though trade-offs are more pronounced in low-resource settings. We argue for evaluation frameworks that integrate efficiency and sustainability alongside accuracy as central dimensions of progress in NLP.
@@ 4.1 Benchmark Evaluation (Flores+)
Despite these reductions in BLEU scores, the distilled model achieved substantial efficiency gains. Inference was 71–78% faster across all three language pairs, with average sentence latency dropping from 0.187s to 0.041s for English-French and from 0.146s to 0.042s for English-Kannada. Carbon emissions per evaluation run decreased by 63–65%, from 0.0075-0.0079 kg to 0.0025-0.0029 kg . These results demonstrate that distillation significantly reduced computational and environmental costs, albeit with modest quality degradation.
| BLEU Score | Inference Time | Carbon Emissions | ||||||
| Language Pair | Full | Distilled | Full | Distilled | Reduction(%) | Full | Distilled | Reduction(%) |
| English-French | 43.99 | 42.10 | 0.187 | 0.041 | 78.1% | 0.0075 | 0.0026 | 65.3% |
| English-Kannada | 15.31 | 11.80 | 0.146 | 0.042 | 71.2% | 0.0079 | 0.0029 | 63.3% |
| English-Hindi | 28.67 | 27.42 | 0.134 | 0.037 | 72.4% | 0.0069 | 0.0025 | 63.8% |
@@ 5 Conclusion
The distilled models reduced inference latency by 71–78% and carbon emissions by 63–65% relative to the full model, while incurring only modest losses in BLEU scores. Performance degradation was most pronounced for low-resource Kannada, whereas high-resource French remained largely stable.
@@
These results demonstrate that distillation and quantization can substantially cut computational demands and environmental costs without severely compromising translation quality. However, the sharper trade-offs observed in low-resource languages highlight the need for careful evaluation before deployment in linguistically diverse settings.
Our analysis focused on inference-time efficiency; the training phase, which contributes significantly to the total carbon footprint of large models, remains an important direction for future research. Expanding to a broader set of languages and tasks will further illuminate multilingual trade-offs and advance the agenda of responsible and environmentally conscious NLP.